# Monotonic normalized heat diffusion for distance-regular graphs with classical parameters of diameter $3$

### v1

@@ 1 Introduction
Due to (1), a necessary condition for the MNHD property is that $r_{t}(u,v)\leq 1$ holds for any $t\geq 0$. However, there do exist graphs for which the function $r_{t}(u,v)$ exceeds 1 for certain vertices $u,v$ at certain time $t$. This is observed by Regev and Shinkar [12, Proposition A.1]. Regev and Shinkar mentioned that this observation was brought to their attention by Cheeger. They also provided an explicit example which is a $4$-regular graph with $10$ vertices [12, Figure 2].
A sufficient condition to ensure that $r_{t}(u,v)\leq 1$ is a heat homogeneity, that is, the heat kernel has constant diagonal, which means that the value of $H_{t}(u,u)$ is independent of the choice of $u$ (see [12, Section 1] and [10, Section 1]). In fact, the heat homogeneity holds if and only if the graph is walk-regular [10, Theorem A.1]. Recall that a graph is walk-regular if for every $\ell\geq 2$, the number of closed walks of length $\ell$ starting from and ending at every vertex is constant. Typical walk-regular graphs includes vertex-transitive graphs, distance-regular graphs, and regular graphs with at most $4$ distinct eigenvalues.
Peres asked in 2013 (see [12, Section 1]) whether the MNHD property holds for all vertex transitive graphs or not. Regev and Shinkar [12] showed that there exists a Cayley graph for which the MNHD property fails. This provides a negative answer to Peres’ question. On the other hand, Price [11] proved that the MNHD property does hold for all finite abelian Cayley graphs.
@@ 2.1 Graph Laplacian and heat kernels
Let $d\geq 1$ be an integer and $G=(V,E)$ a simple finite connected graph with vertex set $V$ and edge set $E$. Suppose that $|V|=n\geq 2$. The graph Laplacian $L$ is a symmetric $n\times n$ matrix $L=(L(x,y))_{x,y\in V}$ defined by $L=\mathrm{Deg}-A$, where $\mathrm{Deg}$ is the diagonal vertex degree matrix and $A$ is the adjacency matrix. The constant vector $\mathbf{1}=(1,1,...,1)^{T}$ is an eigenvector of $L$ corresponding to the eigenvalue 00. Since $G$ is connected, the eigenvalue 00 is simple by the Perron-Frobenius theorem. Let us denote by $\mathrm{sp}(L)$ the set of all distinct eigenvalues of $L$. The graph Laplacian has the following spectral decomposition
@@
[…]
@@ 2.3 Distance-regular graph with classical parameters
A connected graph $G=(V,E)$ of diameter $D$ is distance-regular if and only if it is regular of degree $d$, and if there exist constants $c_{i},a_{i},b_{i}$, known as intersection numbers, such that for all $i=0,1,\dots,D$, and for any two vertices $x,y\in V$ at distance $i=d(x,y)$, among the neighbors of $y$, there are $c_{i}$ at distance $i-1$ from $x$, $a_{i}$ at distance $i$, and $b_{i}$ at distance $i+1$. It follows that $d=b_{0}$, and $a_{i}=d-b_{i}-c_{i}$ for all $i=0,1,\dots,D$. These constants are commonly listed in the so-called intersection array
@@
where $\begin{bmatrix}j\\ 1\\ \end{bmatrix}=1+b+b^{2}+\cdots+b^{j-1}$ is the Gaussian binomial coefficient with basis $b$. To make $b_{i},c_{i}$ feasible parameters, there are necessary restrictions on the numbers $b,\alpha,\beta$. For example, the number $b$ must be an integer not equal to 00 or $-1$ [3, Proposition 6.2.1]. See also Lemma 4 below.
@@ 2.4 Distance-regular graphs with classical parameters of diameter 3
$\begin{array}[]{ll}b_{0}=d=(1+b+b^{2})\beta,&b_{1}=(b+b^{2})(\beta-\alpha),\\ b_{2}=b^{2}(\beta-\alpha(1+b)),&b_{3}=0,\\ c_{0}=0,&c_{1}=1,\\ c_{2}=(1+b)(1+\alpha),&c_{3}=(1+b+b^{2})(1+\alpha(1+b)),\\ a_{0}=0,&a_{1}=-1+b\alpha+b^{2}\alpha+\beta,\\ a_{2}=(1+b)(-1-\alpha+b^{2}\alpha+\beta),&a_{3}=(1+b+b^{2})(-1-\alpha-b\alpha+\beta).\\ \end{array}$ (13)
@@ 3 The MNHD property for regular graphs with four distinct eigenvalues
For any three distinct numbers $\lambda_{i},\lambda_{j},\lambda_{k}\in\mathbb{R}$ with $\{i,j,k\}=\{1,2,3\}$ and any number $L(u,v),L^{2}(u,v)\in\mathbb{R}$, let $\Delta_{i}(u,v),i=1,2,3$ and $\Delta_{ij}(u,v)$ be defined in terms of $\lambda_{i},\lambda_{j},\lambda_{k},L(u,v),L^{2}(u,v)$ as in (7)-(10). Then it holds that
@@
$e^{-t(\lambda_{1}-\lambda_{3})}G^{\prime}(t)=\geq$ $\frac{\lambda_{1}\lambda_{2}}{n}\Delta_{1}(u,v)+\frac{\lambda_{2}\lambda_{1}}{n}\Delta_{2}(u,v)+\frac{\lambda_{3}(\lambda_{1}+\lambda_{2}-\lambda_{3})}{n}\Delta_{3}(u,v)$
@@
Observe from $L=dI-A$ that
@@
=$=\begin{cases}d^{2}+d,&\text{if $x=y$,}\\ -2dA(x,y)+\sum_{w\neq x,y}A(x,w)A(w,y),&\text{if $x\neq y$.}\end{cases}$ (20)
@@ 4 Proof of Theorem 1
$L^{2}\left(x,y\right)=\begin{cases}d^{2}+d,&\text{if}\;d\left(x,y\right)=0;\\ -2d+a_{1},&\text{if}\;d\left(x,y\right)=1;\\ c_{2},&\text{if}\;d\left(x,y\right)=2;\\ 0,&\text{if}\;d\left(x,y\right)\geq 3.\\ \end{cases}$ (21)
@@
Recall that $b$ is an integer not equal to 00 or $-1$. Applying $\beta\geq 1+(1+b)\alpha\geq 1$ from Lemma 4, we deduce that
@@ 4.1 Proof of Theorem 1 in case $b\geq 1$
Inserting into to (33), we obtain
@@
$=$ $b\beta\left\{b^{2}-(1+b+b^{2})(\beta-\alpha)\frac{b_{2}}{c_{3}}.\right\}$
@@
Since $-(1+b+b^{2})(1+\alpha)(1+\alpha+b\alpha)<0$, and the value $1+(1+b)\alpha$ lies in between $\alpha$ and $b+b\alpha+\alpha+1$, we have $g(1+(1+b)\alpha)<0$. Moreover, we check
@@ 4.2 Proof of Theorem 1 in case $b\leq-2$
Using the calculations in the proof of Lemma 8, we check $(i)-(iii)$ as follows.
@@
Known distance regular graphs with classical parameters of diameter at least $3$ are listed in [3, Tables 6.1 and 6.2]. For classification results of distance regular graphs with classical parameters of diameter at least $4$ and $b\leq-2$ or $b=1$, we refer to [3, 6, 13, 14] and the references therein. It is natural to expect the monotonic normalized heat diffusion property holds for any distance regular graphs with classical parameters.
@@ 5 Further Discussion
Moreover, we have $L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})\geq 0$ for any two adjacent vertices $u$ and $v$. Recall that $L^{(}u,v)=-2d+a_{1}=-2d+(d-1-m\gamma)$ and $L(u,v)=-1$. We get $L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})=\frac{1}{2}(1+d+\gamma-m\gamma+M)\geq 0$, as desired.
@@
Case $(iii)$: $d(u,v)=2$. Similar as in Case (i), we have
@@ Acknowledgement
H.Z. is very grateful to Kaizhe Chen and Chenhui Lv for their helpful discussions. This work is supported by National Key R & D Program of China 2023YFA1010200 and National Natural Science Foundation of China No. 12031017 and No. 12431004.

### v2

@@ 1 Introduction
Due to (1), a necessary condition for the MNHD property is that $r_{t}(u,v)\leq 1$ holds for any $t\geq 0$. However, there do exist graphs for which the function $r_{t}(u,v)$ exceeds 1 for certain vertices $u,v$ at certain time $t$. This is observed by Regev and Shinkar [15, Proposition A.1]. Regev and Shinkar mentioned that this observation was brought to their attention by Cheeger. They also provided an explicit example which is a $4$-regular graph with $10$ vertices [15, Figure 2].
A sufficient condition to ensure that $r_{t}(u,v)\leq 1$ is a heat homogeneity, that is, the heat kernel has constant diagonal, which means that the value of $H_{t}(u,u)$ is independent of the choice of $u$ (see [15, Section 1] and [10, Section 1]). In fact, the heat homogeneity holds if and only if the graph is walk-regular [10, Theorem A.1]. Recall that a graph is walk-regular if for every $\ell\geq 2$, the number of closed walks of length $\ell$ starting from and ending at every vertex is constant. Typical walk-regular graphs includes vertex-transitive graphs, distance-regular graphs, and regular graphs with at most $4$ distinct eigenvalues.
Peres asked in 2013 (see [15, Section 1]) whether the MNHD property holds for all vertex transitive graphs or not. Regev and Shinkar [15] showed that there exists a Cayley graph for which the MNHD property fails. This provides a negative answer to Peres’ question. On the other hand, Price [13] proved that the MNHD property does hold for all finite abelian Cayley graphs.
@@
Relevance of the MNHD property in other areas
It is very natural to study whether the MNHD property holds on a given graph. In general, it is highly concerned to understand how certain properties of solutions to differential/difference equations depend on the structure of the underlying graphs. Moreover, the MNHD property is very useful and relevant to other research areas.
It is also very interesting to study the MNHD property in the setting of Riemannian geometry. Similarly, the MNHD property does not always hold for Riemannian manifolds [13, Appendix A]. In the other direction, it was shown that the MNHD property holds on all flat tori [15] and some other Riemannian symmetric spaces including the $3$-dimensional hyperbolic space [13, Section 5 and Appendix B]. It is still open whether the MNHD property holds on all Riemannian symmetric spaces of compact type and on all hyperbolic spaces.
In a lattice in $\mathbb{R}^{n}$, the MNHD property reduces to a monotonicity property of the periodic Gaussian function. An extension of this monotonicity property to the non-spherical Gaussian case [16, Proposition 4.2] can be used to derive a monotonicity property of the normalized heat diffusion on a torus with respect to different flat Riemannian metrics [12, 16]. Such an application also leads to a key ingredient in the framework of the so-called numerical cohomology theory [11, 14].
@@ 2.1 Graph Laplacian and heat kernels
Let $d\geq 1$ be an integer and $G=(V,E)$ a simple finite connected graph with vertex set $V$ and edge set $E$. Suppose that $|V|=n\geq 2$. The graph Laplacian $L$ is a symmetric $n\times n$ matrix $L=(L(x,y))_{x,y\in V}$ defined by $L=\mathrm{Deg}-A$, where $\mathrm{Deg}$ is the diagonal vertex degree matrix and $A$ is the adjacency matrix. The constant vector $\mathbf{1}=(1,1,...,1)^{T}$ is an eigenvector of $L$ corresponding to the eigenvalue 0. Since $G$ is connected, the eigenvalue 0 is simple by the Perron-Frobenius theorem. Let us denote by $\mathrm{sp}(L)$ the set of all distinct eigenvalues of $L$. The graph Laplacian has the following spectral decomposition
@@
where $P_{\lambda}$ is the projection operator onto the eigenspace associated with $\lambda$. The projection operator $P_{0}=\langle\cdot,\frac{\mathbf{1}}{\sqrt{n}}\rangle\frac{\mathbf{1}}{\sqrt{n}}$ corresponding to the simple eigenvalue 0 is an averaging operator, where we use $\langle\cdot,\cdot\rangle$ for the standard Euclidean inner product. In matrix form, we have $P_{0}=\frac{1}{n}J$, with $J$ being the $n\times n$ all-11 matrix.
@@ 2.3 Distance-regular graph with classical parameters
A connected graph $G=(V,E)$ of diameter $D$ is distance-regular if and only if it is regular of degree $d$, and if there exist constants $c_{i},a_{i},b_{i}$, known as intersection numbers, such that for all $i=0,1,\dots,D$, and for any two vertices $x,y\in V$ at distance $i=d(x,y)$, among the neighbors of $y$, there are $c_{i}$ at distance i1i-1 from $x$, $a_{i}$ at distance $i$, and $b_{i}$ at distance $i+1$. It follows that $d=b_{0}$, and $a_{i}=d-b_{i}-c_{i}$ for all $i=0,1,\dots,D$. These constants are commonly listed in the so-called intersection array
@@
where $\begin{bmatrix}j\\ 1\\ \end{bmatrix}=1+b+b^{2}+\cdots+b^{j-1}$ is the Gaussian binomial coefficient with basis $b$. To make $b_{i},c_{i}$ feasible parameters, there are necessary restrictions on the numbers $b,\alpha,\beta$. For example, the number $b$ must be an integer not equal to 0 or 1-1 [3, Proposition 6.2.1]. See also Lemma 4 below.
@@ 2.4 Distance-regular graphs with classical parameters of diameter 3
[…]
@@ 3 The MNHD property for regular graphs with four distinct eigenvalues
For any three distinct numbers $\lambda_{i},\lambda_{j},\lambda_{k}\in\mathbb{R}$ with $\{i,j,k\}=\{1,2,3\}$ and any numbers $L(u,v),L^{2}(u,v)\in\mathbb{R}$, let $\Delta_{i}(u,v),i=1,2,3$ and $\Delta_{ij}(u,v)$ be defined in terms of $\lambda_{i},\lambda_{j},\lambda_{k},L(u,v),L^{2}(u,v)$ as in (7)-(10). Then it holds that
@@
$e^{-t(\lambda_{1}-\lambda_{3})}G^{\prime}(t)\geq$ $\frac{\lambda_{1}\lambda_{2}}{n}\Delta_{1}(u,v)+\frac{\lambda_{2}\lambda_{1}}{n}\Delta_{2}(u,v)+\frac{\lambda_{3}(\lambda_{1}+\lambda_{2}-\lambda_{3})}{n}\Delta_{3}(u,v)$
@@
Observe from L=dIAL=dI-A that
@@
$=\begin{cases}d^{2}+d,&\text{if $x=y$,}\\ -2dA(x,y)+\sum_{w\neq x,y}A(x,w)A(w,y),&\text{if $x\neq y$.}\end{cases}$ (20)
@@ 4 Proof of Theorem 1
[…]
@@
Recall that $b$ is an integer not equal to 0 or 1-1. Applying $\beta\geq 1+(1+b)\alpha\geq 1$ from Lemma 4, we deduce that
@@ 4.1 Proof of Theorem 1 in case $b\geq 1$
Inserting into (33), we obtain
@@
$=$ $b\beta\left\{b^{2}-(1+b+b^{2})(\beta-\alpha)\frac{b_{2}}{c_{3}}\right\}.$
@@
Since $(1+b+b^{2})(1+\alpha)(1+\alpha+b\alpha)<0$, and the value $1+(1+b)\alpha$ lies in between $\alpha$ and $b+b\alpha+\alpha+1$, we have $g(1+(1+b)\alpha)<0$. Moreover, we check
@@ 4.2 Proof of Theorem 1 in case $b\leq-2$
Using the calculations in the proof of Lemma 8, we check (i)(iii)(i)-(iii) as follows.
@@
Known distance regular graphs with classical parameters of diameter at least $3$ are listed in [3, Tables 6.1 and 6.2]. For classification results of distance regular graphs with classical parameters of diameter at least $4$ and $b\leq-2$ or $b=1$, we refer to [3, 6, 17, 18] and the references therein. It is natural to expect the monotonic normalized heat diffusion property holds for any distance regular graphs with classical parameters. Recall that Proposition 1 provides explicit expressions for all $D+1$ distinct Laplacian eigenvalues for any distance-regular graph with classical parameters $(D,b,\alpha,\beta)$. With this proposition in hand, the approach developed herein may pave the way for confirming the MNHD property for the cases where the number of distinct eigenvalues is greater than $4$.
@@ 5 Further Discussion
Moreover, we have $L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})\geq 0$ for any two adjacent vertices $u$ and $v$. Recall that $L(u,v)=-2d+a_{1}=-2d+(d-1-m\gamma)$ and L(u,v)=1L(u,v)=-1. We get $L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})=\frac{1}{2}(1+d+\gamma-m\gamma+M)\geq 0$, as desired.
@@
Case $(iii)$: $d(u,v)=3$. Similar as in Case (i), we have
@@ Acknowledgement
We are very grateful to the anonymous referee for suggestions that improved the readability of our manuscript. H.Z. is very grateful to Kaizhe Chen and Chenhui Lv for their helpful discussions. This work is supported by National Key R & D Program of China 2023YFA1010200 and National Natural Science Foundation of China No. 12031017 and No. 12431004.
Availability of Data
No data was used for the research described in the article.
Declarations of Conflict of Interest
The authors do not have any possible conflicts of interest.